Two numbers have a difference of 10 and the sum of 34. What are the numbers?
Answer by checkley77(12844)

x+y=84 or x=84-y
y-x=10
y-(84-y)=10
y-84+y=10
2y=10+84
y=94/2
y=47 answer
47-x=10
-x=10-47
-x=-37
x=37 answer.
Proof:
37+47=84
84=84